OpenStudy (anonymous):

The length of rectangle is 23 cm less than 6 times its width. It's area is 35 square cm. Find Dimensios. Show me how to do this please? :)

OpenStudy (anonymous):

*dimensions sorry :D

OpenStudy (anonymous):

l+23=6w A=lw=35

OpenStudy (anonymous):

w(6w-23)=35 6w^2-23w-35=0

OpenStudy (anonymous):

6w^2-30w+7w-35=0 6w(w-5)+7(w-5)=0 (w-5)(6w+7)=0 w=5 or -7/6 We can discard -7/6 because the dimensions will not be negative. l=A/w=35/5=7

OpenStudy (anonymous):

7+23=6(5) so it looks good